# Agentar-Fin-R1: Enhancing Financial Intelligence through Domain Expertise, Training Efficiency, and Advanced Reasoning

### v3

@@ Abstract
Large Language Models (LLMs) exhibit considerable promise in financial applications; however, prevailing models frequently demonstrate limitations when confronted with scenarios that necessitate sophisticated reasoning capabilities, stringent trustworthiness criteria, and efficient adaptation to domain-specific requirements. We introduce the Agentar-Fin-R1 series of financial large language models (8B and 32B parameters), specifically engineered based on the Qwen3 foundation model to enhance reasoning capabilities, reliability, and domain specialization for financial applications. Our optimization approach integrates a high-quality, systematic financial task label system with a comprehensive multi-layered trustworthiness assurance framework. This framework encompasses high-quality trustworthy knowledge engineering, multi-agent trustworthy data synthesis, and rigorous data validation governance. Through label-guided automated difficulty-aware optimization, tow-stage training pipeline, and dynamic attribution systems, we achieve substantial improvements in training efficiency. Our models undergo comprehensive evaluation on mainstream financial benchmarks including Fineva, FinEval, and FinanceIQ, as well as general reasoning datasets such as MATH-500 and GPQA-diamond. To thoroughly assess real-world deployment capabilities, we innovatively propose the Finova evaluation benchmark, which focuses on agent-level financial reasoning and compliance verification. Experimental results demonstrate that Agentar-Fin-R1 not only achieves state-of-the-art performance on financial tasks but also exhibits exceptional general reasoning capabilities, validating its effectiveness as a trustworthy solution for high-stakes financial applications. The Finova bench is available at https://github.com/antgroup/Finova.
@@ 1 Introduction
Figure 1: Agentar-Fin-R1-32B performance on financial benchmarks. Results on general reasoning benchmark (MATH-500, GPQA-diamond) are detailed in following sections.
@@
Our experimental results demonstrate that Agentar-Fin-R1 achieves superior performance across financial benchmarks (Fineva, FinEval, FinanceIQ and Finova) while maintaining competitive results on general reasoning tasks (MATH-500, GPQA-diamond), validating the effectiveness of domain-specialized optimization without catastrophic forgetting.
@@ 2.3.3 Verification and Checking: Rigorous Multi-Modal Validation
where $\text{sim}(\cdot,\cdot)$ represents a semantic similarity function that incorporates both lexical overlap and contextual embedding similarity measures to capture nuanced agreement patterns.
@@ 4.1 Benchmark
  • Fineva[1]: A financial benchmark encompassing 33 sub-dimensions across five core capability categories: financial cognition, financial knowledge, financial logic, content generation, and security & compliance.
  • FinEval[12]: A benchmark focused on evaluating financial question-answering models, covering a variety of financial topics and scenarios.
@@ 4.2 Training detail
We train Agentar-Fin-R1-8B and Agentar-Fin-R1-32B based on Qwen3-8B-Instruct and Qwen3-32B-Instruct respectively. The training process consists of two stages: initial SFT, and then GRPO and SFT refinement. For the 8B model, we use 16 NVIDIA A100 GPUs, while the 32B model is trained on 64 A100 GPUs. All training uses bf16 precision with a sequence length of 16K and appropriate gradient accumulation steps to ensure training stability. Beyond the synthetically generated data derived from our data synthesis framework detailed in Section 2, our dataset incorporates financial reasoning data from our proprietary Agentar-DeepFinance-300K[32], general-purpose training corpora[23], as well as datasets sourced from Llama-Nemotron[2] and openthoughts[10].
@@ 4.4 Main Results
Fineva FinEval FinanceIQ Finova Avg. MATH GPQA Avg. Avg.
@@
Qwen2.5-14B-Instruct 14B 84.41 71.60 68.82 37.95 65.70 79.40 35.35 57.38 62.92
Qwen2.5-72B-Instruct 72B 87.14 76.64 74.03 48.22 71.51 82.60 43.43 63.02 68.68
DeepSeek-V3 671B 87.69 77.99 73.93 54.29 73.48 88.40 52.53 70.47 72.47
GPT-4o - 85.82 74.26 72.18 45.20 69.37 78.80 51.01 64.91 67.88
@@
Qwen3-8B 8B 84.36 76.27 73.06 54.45 72.04 93.80 59.60 76.70 73.59
Qwen3-32B 32B 88.94 80.50 78.03 59.46 76.73 95.40 63.13 79.27 77.58
Qwen-QwQ-32B 32B 89.63 82.69 81.58 61.70 78.90 93.60 61.62 77.61 78.47
DeepSeek-R1 671B 88.04 84.93 83.98 61.28 79.56 95.20 72.39 83.80 80.97
GPT-o1 - 88.28 81.32 78.72 60.46 77.20 94.80 76.77 85.79 80.06
@@
Qwen-Fin-R1-7B 7B 79.18 65.80 61.6 38.37 61.24 72.80 28.28 50.54 57.67
Qwen-Dianjin-R1-7B 7B 84.30 74.90 72.89 41.89 68.50 74.60 41.92 58.26 65.08
Qwen-Dianjin-R1-32B 32B 88.47 80.41 81.49 56.02 76.60 84.40 58.59 71.50 74.90
XuanYuan-FinX1-Preview 70B 81.85 69.03 70.02 50.74 67.91 71.80 42.42 57.11 64.31
Agentar-Fin-R1-8B 8B 91.52 85.09 84.24 63.56 81.10 93.40 60.10 76.75 79.65
Agentar-Fin-R1-32B 32B 92.38 87.70 86.79 69.93 84.20 93.80 68.18 80.99 83.13
Table 2: Performance comparison in accuracy across financial benchmarks (Fineva, FinEval, FinanceIQ, Finova) and general reasoning tasks (MATH: MATH-500, GPQA: GPQA-diamond). Scores in bold indicate the best results. Results include individual benchmark performance and averaged scores for financial tasks (Financial Avg.), general reasoning (General Avg.), and overall performance (Overall Avg.). Agentar-Fin-R1-32B achieves state-of-the-art performance across all financial benchmarks, as well as competitive results on general reasoning.
Figure 5: Performance comparison between Agentar-Fin-R1 and Qwen3 models (8B and 32B variants) on financial benchmarks (Fineva, FinEval, FinanceIQ, Finova) and general reasoning benchmarks (MATH: MATH-500, GPQA: GPQA-diamond). Our proposed models show improvements across both domain-specific and general reasoning tasks.
The comprehensive evaluation results presented in Table 2 demonstrate the superior performance of our Agentar-Fin-R1 models across both financial and general domains. Our Agentar-Fin-R1-32B model establishes a new state-of-the-art benchmark with an average score of 83.13, substantially surpassing all competing baselines. Particularly noteworthy is the consistent dominance of our models across the entire spectrum of financial evaluation benchmarks: Agentar-Fin-R1-32B achieves optimal performance on Fineva (92.38), FinEval (87.70), FinanceIQ (86.79), and Finova (69.93), while the more parameter-efficient Agentar-Fin-R1-8B variant maintains highly competitive performance despite its reduced computational footprint.
@@ 4.5.1 Ablation Study on Label System and Weighted Training Framework
All configurations are evaluated on the same downstream tasks, using the benchmark datasets: Fineva, FinEval, FinanceIQ, Finova, MATH-500, and GPQA-diamond. We report task-specific accuracy and overall average performance. The model used for the experiment is Qwen3-8B.
@@
Fineva FinEval FinanceIQ Finova MATH GPQA All Datasets
Random Sampling (50% data) 86.43 79.23 76.72 58.73 92.20 58.59 75.32
Label-Only Stratified Sampling (50% data) 88.61 82.98 78.43 61.32 92.00 57.07 76.74
Ours (10% data) 88.14 81.94 77.22 61.01 93.20 58.59 76.68
Ours (30% data) 89.34 83.46 78.13 61.28 91.80 60.10 77.35
Ours (50% data) 89.97 84.24 79.91 62.92 92.60 60.10 78.12
Full Data (Vanilla SFT) 89.23 83.89 78.69 61.63 91.80 58.08 77.22
Base Model (Qwen3-8B) 84.36 76.27 73.06 54.45 93.80 59.60 73.59
@@
  • The performance progressively improves as we increase the data budget: 10% → 30% → 50% (76.68 → 77.35 → 78.12).
@@
  • Label System Impact: Label-only stratified sampling (76.74) outperforms random sampling (75.32) by 1.42, demonstrating the value of structured sampling.
  • Weighting Mechanism Impact: Our full framework (78.12) further improves upon label-only sampling by 1.38, validating the effectiveness of difficulty-aware weighting.
  • Combined Effect: The synergy between label system and weighting mechanism provides a total improvement of 2.80 over random sampling.
@@ 4.5.2 Ablation Study on Two-Stage Training Pipeline
All configurations are evaluated on the same benchmark datasets: Fineva, FinEval, FinanceIQ, Finova, MATH-500, and GPQA-diamond.
@@
Fineva FinEval FinanceIQ Finova MATH GPQA All Datasets
Single-Stage (SFT Only) 90.21 84.19 82.32 62.87 94.20 58.59 78.73
Ours (Two-Stage) 91.52 85.09 84.24 63.56 93.40 59.60 79.57
Base Model (Qwen3-8B) 84.36 76.27 73.06 54.45 93.80 59.60 73.59

### v4

@@ Abstract
Large Language Models (LLMs) exhibit considerable promise in financial applications; however, prevailing models frequently demonstrate limitations when confronted with scenarios that necessitate sophisticated reasoning capabilities, stringent trustworthiness criteria, and efficient adaptation to domain-specific requirements. We introduce the Agentar-Fin-R1 series of financial large language models (8B and 32B parameters), specifically engineered based on the Qwen3 foundation model to enhance reasoning capabilities, reliability, and domain specialization for financial applications. Our optimization approach integrates a high-quality, systematic financial task label system with a comprehensive multi-layered trustworthiness assurance framework. This framework encompasses high-quality trustworthy knowledge engineering, multi-agent trustworthy data synthesis, and rigorous data validation governance. Through label-guided automated difficulty-aware optimization, tow-stage training pipeline, and dynamic attribution systems, we achieve substantial improvements in training efficiency. Our models undergo comprehensive evaluation on mainstream financial benchmarks including FinEval 1.0, and FinanceIQ, as well as general reasoning datasets such as MATH-500 and GPQA-diamond. To thoroughly assess real-world deployment capabilities, we innovatively propose the Finova evaluation benchmark, which focuses on agent-level financial reasoning and compliance verification. Experimental results demonstrate that Agentar-Fin-R1 not only achieves state-of-the-art performance on financial tasks but also exhibits exceptional general reasoning capabilities, validating its effectiveness as a trustworthy solution for high-stakes financial applications. The Finova bench is available at https://github.com/antgroup/Finova.
@@ 1 Introduction
Figure 1: Agentar-Fin-R1-32B performance: significantly outperforms on financial benchmarks (FinEval 1.0, FinanceIQ, Finova), and on general reasoning benchmarks (MATH-500, GPQA-diamond), outperforms comparable-sized general models while achieving performance comparable to large-scale models like DeepSeek R1 and GPT-o1.
@@
Our experimental results demonstrate that Agentar-Fin-R1 achieves superior performance across financial benchmarks (FinEval 1.0, FinanceIQ and Finova) while maintaining competitive results on general reasoning tasks (MATH-500, GPQA-diamond), validating the effectiveness of domain-specialized optimization without catastrophic forgetting.
@@ 2.3.3 Verification and Checking: Rigorous Multi-Modal Validation
[…]
@@ 4.1 Benchmark
  • FinEval 1.0[11]: A benchmark focused on evaluating financial question-answering models, covering a variety of financial topics and scenarios.
@@ 4.2 Training detail
We train Agentar-Fin-R1-8B and Agentar-Fin-R1-32B based on Qwen3-8B-Instruct and Qwen3-32B-Instruct respectively. The training process consists of two stages: initial SFT, and then GRPO and SFT refinement. For the 8B model, we use 16 NVIDIA A100 GPUs, while the 32B model is trained on 64 NVIDIA A100 GPUs. All training uses bf16 precision with a sequence length of 16K and appropriate gradient accumulation steps to ensure training stability. Beyond the synthetically generated data derived from our data synthesis framework detailed in Section 2, our dataset incorporates financial reasoning data from our proprietary Agentar-DeepFinance-100K[31], general-purpose training corpora[22], as well as datasets sourced from Llama-Nemotron[1] and openthoughts[9].
@@ 4.4 Main Results
FinEval 1.0 FinanceIQ Finova Avg. MATH GPQA Avg. Avg.
@@
Qwen2.5-14B-Instruct 14B 71.60 68.82 37.95 59.46 79.40 35.35 57.38 58.62
Qwen2.5-72B-Instruct 72B 76.64 74.03 48.22 66.30 82.60 43.43 63.02 64.98
DeepSeek-V3 671B 77.99 73.93 54.29 68.74 88.40 52.53 70.47 69.43
GPT-4o - 74.26 72.18 45.20 63.88 78.80 51.01 64.91 64.29
@@
Qwen3-8B 8B 76.27 73.06 54.45 67.93 93.80 59.60 76.70 71.44
Qwen3-32B 32B 80.50 78.03 59.46 72.66 95.40 63.13 79.27 75.30
Qwen-QwQ-32B 32B 82.69 81.58 61.70 75.32 93.60 61.62 77.61 76.24
DeepSeek-R1 671B 84.93 83.98 61.28 76.73 95.20 72.39 83.80 79.56
GPT-o1 - 81.32 78.72 60.46 73.50 94.80 76.77 85.79 78.25
@@
Qwen-Fin-R1-7B 7B 65.80 61.6 38.37 55.26 72.80 28.28 50.54 53.37
Qwen-Dianjin-R1-7B 7B 74.90 72.89 41.89 63.23 74.60 41.92 58.26 61.24
Qwen-Dianjin-R1-32B 32B 80.41 81.49 56.02 72.64 84.40 58.59 71.50 72.18
XuanYuan-FinX1-Preview 70B 69.03 70.02 50.74 63.26 71.80 42.42 57.11 60.80
Agentar-Fin-R1-8B 8B 85.09 84.24 63.56 77.63 93.40 60.10 76.75 77.41
Agentar-Fin-R1-32B 32B 87.70 86.79 69.93 81.47 93.80 68.18 80.99 81.28
Table 2: Performance comparison in accuracy across financial benchmarks (FinEval 1.0, FinanceIQ, Finova) and general reasoning tasks (MATH: MATH-500, GPQA: GPQA-diamond). Scores in bold indicate the best results. Results include individual benchmark performance and averaged scores for financial tasks (Financial Avg.), general reasoning (General Avg.), and overall performance (Overall Avg.). Agentar-Fin-R1-32B achieves state-of-the-art performance across all financial benchmarks, as well as competitive results on general reasoning.
Figure 5: Performance comparison between Agentar-Fin-R1 and Qwen3 models (8B and 32B variants) on financial benchmarks (FinEval 1.0, FinanceIQ, Finova) and general reasoning benchmarks (MATH: MATH-500, GPQA: GPQA-diamond). Our proposed models show improvements across both domain-specific and general reasoning tasks.
The comprehensive evaluation results presented in Table 2 demonstrate the superior performance of our Agentar-Fin-R1 models across both financial and general domains. Our Agentar-Fin-R1-32B model establishes a new state-of-the-art benchmark with an average score of 81.28, substantially surpassing all competing baselines. Particularly noteworthy is the consistent dominance of our models across the entire spectrum of financial evaluation benchmarks: Agentar-Fin-R1-32B achieves optimal performance on FinEval 1.0 (87.70), FinanceIQ (86.79), and Finova (69.93), while the more parameter-efficient Agentar-Fin-R1-8B variant maintains highly competitive performance despite its reduced computational footprint.
@@ 4.5.1 Ablation Study on Label System and Weighted Training Framework
All configurations are evaluated on the same downstream tasks, using the benchmark datasets: FinEval 1.0, FinanceIQ, Finova, MATH-500, and GPQA-diamond. We report task-specific accuracy and overall average performance. The model used for the experiment is Qwen3-8B.
@@
FinEval 1.0 FinanceIQ Finova MATH GPQA All Datasets
Random Sampling (50% data) 79.23 76.72 58.73 92.20 58.59 73.09
Label-Only Stratified Sampling (50% data) 82.98 78.43 61.32 92.00 57.07 74.36
Ours (10% data) 81.94 77.22 61.01 93.20 58.59 74.39
Ours (30% data) 83.46 78.13 61.28 91.80 60.10 74.75
Ours (50% data) 84.24 79.91 62.92 92.60 60.10 75.95
Full Data (Vanilla SFT) 83.89 78.69 61.63 91.80 58.08 74.82
Base Model (Qwen3-8B) 76.27 73.06 54.45 93.80 59.60 71.44
@@
  • The performance progressively improves as we increase the data budget: 10% → 30% → 50% (74.39 → 74.75 → 75.95).
@@
  • Label System Impact: Label-only stratified sampling (74.36) outperforms random sampling (73.09) by 1.27, demonstrating the value of structured sampling.
  • Weighting Mechanism Impact: Our full framework (75.95) further improves upon label-only sampling by 1.59, validating the effectiveness of difficulty-aware weighting.
  • Combined Effect: The synergy between label system and weighting mechanism provides a total improvement of 2.86 over random sampling.
@@ 4.5.2 Ablation Study on Two-Stage Training Pipeline
All configurations are evaluated on the same benchmark datasets: FinEval 1.0, FinanceIQ, Finova, MATH-500, and GPQA-diamond.
@@
FinEval 1.0 FinanceIQ Finova MATH GPQA All Datasets
Single-Stage (SFT Only) 84.19 82.32 62.87 94.20 58.59 76.43
Ours (Two-Stage) 85.09 84.24 63.56 93.40 59.60 77.18
Base Model (Qwen3-8B) 76.27 73.06 54.45 93.80 59.60 71.44